OpenStudy (anonymous):

Solve the following system using the substitution method. x + 3y = -5 5x + 6y = -19 negative three, negative two-thirds negative three, two-thirds three, negative two-thirds three, two-thirds

OpenStudy (tkhunny):

That first x looks compelling. x = -5 - 3y Substitute this value for x into the 2nd equation.

OpenStudy (anonymous):

x + 3y = -5 i.e. x= -3y-5 Substituting x= -3y-5 in 5x + 6y = -19 we find 5(-3y-5) + 6y = -19 i.e. -15y -25 +6y =-19 i.e. 6y-15y =25-19 i.e. -9y =6 i.e. y =6/(-9) i.e. y =2/(-3) Therefore x = -3*2/(-3)-5 x = 2-5 x = -3 thus {x, y} ={-3, 2/(-3)} is the required solution set of the given system of linear equations.

OpenStudy (anonymous):

@acriley3

OpenStudy (anonymous):

So what would the ordered pair be?

OpenStudy (anonymous):

I mean the answer

OpenStudy (anonymous):

Out of the ones I put up there

OpenStudy (anonymous):

B?

OpenStudy (tkhunny):

A and C have y = -2/3 - You must pick between those.